OpenStudy (anonymous):

√49 ---- √2 RADICAL EXPRESSION PLZ DONT ANSWER AS A DECIMAL

OpenStudy (anonymous):

first rationalize denominator.

OpenStudy (anonymous):

multiply numerator and denominator by \[\sqrt{2}\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

\[\sqrt{49}\]=7

OpenStudy (anonymous):

\[7\sqrt{2} \div2\]